2. Two people are stranded on an island where they catch fish and gather coconuts daily:

Let f=fish, and cc=coconuts.

Person 1 can catch 10 fish (f) or gather 10 coconuts (cc) in a day.

Person 2 can catch 2 fish (f), or gather 8 coconuts (cc) in a day.

Which of the following is true?

A. They can both benefit from trade if they trade at 1 (f) to 1 (cc).

B. They can both benefit from trade if they trade at 1 (f) to 4 (cc).

C. They can both benefit from trade if they trade at 1 (f) to 9 (cc).

D. They can both benefit from trade if they trade at 1(F) to (2 or 3)(cc).
Answer #1
Opportunity cost of catching a fish in terms of coconut Opportunity cost of gathering a coconut in terms of fish
Person 1 (10/10) = 1 coconut (10/10) = 1 fish
Person 2 (8 /2) = 4 coconut (2 /8) = 0.25 fish

Opportunity cost of catching a fish is lower for person 1. Therefore, Person 1 has a comparative advantage in the catching a fish.

Similarly, Person 2 has a comparative advantage in gathering a coconut.

Person 1 is ready to exchange one fish with coconut for at least 1 coconut.

Person 2 is ready to accept one fish in exchange of coconut for at most 4 units of coconut.

Exchange rate of fish with coconut agreed by both (i.e., person 1 and person 2)

1 coconut < 1 fish < 4 coconut.

Hence, They can both benefit from trade if they trade at 1 (f) to 4 (cc).

Answer: Option (B).